OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the x's to add out? 2x + 3y = 8 -6x + 4y = 11

razor99 (razor99):

Thats simultaneous right

OpenStudy (anonymous):

addition methods

OpenStudy (anonymous):

might be what u said

razor99 (razor99):

i got the answer lol

razor99 (razor99):

X=-1/26 Y=35/13

OpenStudy (anonymous):

3 -3 6 are my option

OpenStudy (zzr0ck3r):

we want to multiply the first equation by a number so that when we add the two equations the x's cancel out so we want the x in the first equation to be 6x what do we multiply 2x by to make it 6x?

OpenStudy (anonymous):

3?

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

we want it to be 6x so that when we add the first equation to the second the x's "go away" 6x+(-6x)=0

OpenStudy (anonymous):

yayyy what about 3x - y = 3 -2x + 2y = 6

OpenStudy (zzr0ck3r):

well we have -2x on the bottom so we want 2x on the top what do we multiply 3x by to get 2x?

OpenStudy (zzr0ck3r):

\[3x*\frac{?}{?}=2x\]

OpenStudy (anonymous):

answer be 2?

OpenStudy (zzr0ck3r):

2*3x = 2x?

OpenStudy (anonymous):

noo 6x ?

OpenStudy (zzr0ck3r):

\[3x*\frac{2}{3}=\cancel{3}x*\frac{2}{\cancel{3}}=2x\]

OpenStudy (anonymous):

so the answer would be 6?

OpenStudy (zzr0ck3r):

no I asked you what do we need to multiply 3x by to get 2x you answerd 2 I showed you that 2*3x does not equal 3x it equals 6x

OpenStudy (zzr0ck3r):

what did I multiply 3x by to get 2x ?

OpenStudy (anonymous):

im not sure :(